Hello.
Could someone please help me out with the following question thanks so much;
A local fitness club offers two different formulas to pay.

Formula A: monthly payments of 25$ plus another 2$ each time you visit
Formula B: 7$ each time you visit
A) determine the number of visites for each formula to be equal amounts.
B) What formula is cheaper for one month?
Explain your answer. 30 days is the number of months.
The equation i can up with is formula a: c= 2m+ 25
formula b: c=7m

I just cant seem to get it.
If you could please help me out i would greatly appreciate it
-kim
thank you again

Answer by uma(370)

Kim your equations are right.
C = 2m + 25 in formula 1 and C = 7m in formula 2
Equating them we have,
7m = 2m + 25
==> 5m = 25
==> m = 5
Therefore for 5 visits, the amounts are equal.
Cost by formula 1 for 30 days = 2(30) + 25
= 60 + 25 = 85$
Cost by formula 2 for 30 days = 7(30)
= 210$
Therefore formula 1 is cheaper for a month.